alones (alones):

Ughhhh I KEEP ON GETTING -6!!

alones (alones):

What is the value of b2 - 4ac for the following equation? \[2x ^2 - 2x - 1 = 0\]

OpenStudy (erak):

c is "-1", that's important

Directrix (directrix):

Discriminant Study Aid

alones (alones):

Uhh so basically I wrote \[A=2x^2 \] B=2x C=1 And I got them together 2-4(2)(1)

Directrix (directrix):

b² - 4 a * c Do you agree that: a = 2 b = -2 c = -1

alones (alones):

Oh damn -1 i did just 1 /.<

Directrix (directrix):

b² - 4 a * c ( -2 )² - 4 * ( 2) * (-1) =

Directrix (directrix):

( -2 )² = ?

alones (alones):

Oh an -2 too 4 woudl be the asnwer

Directrix (directrix):

Next: 4 * ( 2) * (-1) =

alones (alones):

-8?

Directrix (directrix):

ok

alones (alones):

ok?

Directrix (directrix):

Putting the problem back together: ( -2 )² - 4 * ( 2) * (-1) = 4 - ( -8) =

Directrix (directrix):

4 - ( -8) =

alones (alones):

Sorry this lag is just killing me, okay my answer is 12.

Directrix (directrix):

Correct

Directrix (directrix):

( -2 )² - 4 * ( 2) * (-1) = 12

alones (alones):

Oh thanks<3

Directrix (directrix):

Alrighty, then.